# STORMY : A Real-time Triggering Framework using Yamagawa Solar Spectrograph for Active Solar Emission Observations with the MWA

### v1

@@ 3.3 Background Subtraction
Here, $M_{T}[S^{\prime}(\nu=\nu_{0},t)]$ represents the moving median of pixel values from timestamp tTt-T to $t$ for frequency $\nu_{0}$. To get good estimates of the background noise, we stitch together two consecutive 30-minute data chunks and calculate the median by setting the window width, $T$, to 1800 seconds. This duration is significantly larger than the lifetime of most solar radio bursts at any given frequency, and has been chosen with the motivation of reducing the likelihood of any solar radio bursts present in the window impacting the median value.
@@
Finally, this background-subtracted dynamic spectrum, $S_{\mathrm{bkgsub}}(\nu,t)$, is used for further analysis. The top panels of figure 7 shows the dynamic spectrum before and after the median based background subtraction. As it can be clearly seen from the figure that the solar event has become much prominent.
@@ 4.1 Generating Binary Dynamic Spectrum
The temporal resolution of the data received from the Yamagawa is 1 second and hence, to avoid false detection due to the presence of any low-level transient RFI, we only search for events which are present for more than at least 3 seconds. For each spectral channel, we estimate the background levels by calculating the standard deviation ($\sigma(\nu)$) of from the previous day quiet time dynamic spectrum. A signal-to-noise ratio (SNR) dynamic spectrum of the real-time data is created after the following operation on each spectral channel –
@@ 4.2 Event Detection Algorithm
We try to find closed contours on the image after morphological closing operation using OpenCV which uses border following method suzuki1985topological. Contours having minimum span of 15 MHz in frequency and 3 seconds in time are considered to be true events. This method not only finds closed contours but also filters out residual RFI which do not match the frequency and temporal span criteria. The bottom panels in figure 7 show the detection of such a closed contour region in the binary dynamic spectrum. If we find any such region within the most recent two minutes, we consider this event to be observed by triggering the MWA.
@@ 5 Triggering the MWA
The final step is to communicate with the MWA once a triggering condition is met and to schedule appropriate solar observations. This is achieved via a web-based triggering mechanism that has been developed for the MWA Hancock. On arrival of a trigger, the buffer data is frozen and dumped to disk. As mentioned in Sec. 2.1, the buffer data is in ‘picket-fence’ mode. Simultaneously, an interferometric observation for the next 2 minutes is initiated. STORMY’s detection algorithm estimates the spectral span of the detected event. This information is used for defining the spectral coverage of the triggered observations. If $\nu_{min}$ and $\nu_{max}$ are the minimum and maximum detected frequency, the MWA observing bandwidth of 30.72 MHz is distributed linearly across $(\nu_{min}-30)$ MHz to $(\nu_{max}+10)$ MHz, subject to the MWA observing band of 8080300MHz300\ MHz. The 30MHz30\ MHz guard band on the lower frequency side is chosen to accommodate the typical drift rate of type II bursts of $0.1\ MHz\ s^{-1}$. Figure 8 shows how STORMY chooses the spectral channels for interferometric observations. The 30 MHz in lower limit and 10 MHz in upper limit are guard bands which allow for spectral drifts and the presence of weaker active emissions in the neighboring spectral regions which might not have been detectable at Yamagawa, but might be detectable by the more sensitive MWA.
Figure 8: This figure shows the spectral coverage of buffer channels and how STORMY schedules MWAX observation after detection. The yellow rectangles are buffer channels. When STORMY triggers an observation (denoted by blue vertical line), the 30.72 MHz band width is allocated according to the detected spectral span (red rectangles).
@@ 6 Quantifying the performance of STORMY
Figure 10: Top panels: (Left) Pre-processed Yamagawa dynamic spectrum. (Right) Radio image of the data captured at 126 MHz on 1 August 2024.The dotted white circle shows the optical disc of the Sun. The red contours are at the 1,10,30,50,90% of the peak. The ellipse inside the box at the bottom left corner shows the point-spread-function. Bottom panels: (Left and Middle) close-up of AIA images of during the event. The dotted circles show the active region from which the eruption is taking place. (Right) Radio contours overlaid on the AIA map.
@@ 7 Some Sample Events Captured by STORMY
[…]
8 Discussion and Conclusion
Though developed with the objective of increasing the observing efficiency of solar activity by the MWA, the framework of STORMY has wider applicability. The following text discusses a few specific instances of where solar triggering can be implemented on other modern radio interferometers:
  • LOFAR: The LOw Frequency ARrray is a versatile instrument, primarily due to its fully digital signal chain. It can observe multiple sources simultaneously via its multi-beam and subarray capabilities. The upcoming LOFAR 2.0 (LOFAR 2.0 white paper) includes a LOFAR4SW (LOFAR for Space Weather) extension, which will use few stations to track the Sun while it is above the instruments elevation limit commensal with other astronomical observations for both the high band antennas (HBA) and low band antennas (LBA). Using a part of the same instrument for solar monitoring offers several advantages over using a remote spectrograph - (a) the trigger comes from a monitoring system of comparable sensitivity and experiences RFI environment identical to the observing system, (b) co-location of the monitoring station removes the dependency on an independent distant system and eliminates the delays arising from the time taken to fetch remote data, substantially reducing the storage requirement for the ring buffer. Additionally, in the case of LOFAR, having monitoring observations in HBA can potentially enable one to raise an observing trigger for the LBA just a little before the emission drifts into that observing band. As an illustration, we demonstrate an application to LOFAR data in figure 12. These data come from the LBA beam-formed mode with spectral resolution of 160 kHz and temporal resolution of 1 s. STORMY’s RFI excision and burst detection algorithms are seen to work very well on the LOFAR dynamic spectra. Hence, STORMY can be used on real-time beam formed data to detect events and trigger interferometric observation when appropriate.
@@ 8 Discussion and Conclusion
  • SKAO: The triggering aspects are particularly relevant for the SKA-Low, which operates in the 50–350 MHz band where the bulk of solar radio activity is observed. Like the MWA, it uses aperture array elements and hence electronic pointing, essentially eliminating the slew time needed to point to the target source, post the trigger. The SKA-Low is expected to reach its AA* milestone before the end of this decade SKAO-YearInLife. It will have 307 stations, each comprising 256 dual-polarization dipoles log-periodic antennas placed over a footprint of 38m38\ m diameter. The SKA-Low supports flexible options for forming not only subarrays, but also substations, with the latter providing access to larger FoV as well as shorter baselines (within a station), though subject to bandwidth restrictions imposed by the correlator and the network capacity. AA* will support up to 1440 substations. The smallest substations have a 6m6\ m foot print and typically include 6 dual-pol antennas SKAO-Low-SubStation-Templates. A subarray comprising one such substation from $\sim$60 SKA-Low stations within a diameter of $\sim 10\ km$ will be an excellent match to the needs of triggered solar observing and is a referred to as solar subarray in the following text. It uses <0.5<0.5% of the SKA-Low collecting area, and hence has minimal sensitivity impact on the observations it is commensal with. Along the lines of the picket fence mode routinely employed at the MWA for solar observations, the solar subarray can use narrow spectral channels distributed across the 50–350 MHz SKA-Low observing band. For instance, using a hundred spectral channels, each 5.4 kHz in bandwidth, amounts to a mere 0.54 MHz of bandwidth. The solar subarray can be used to fill the voltage buffer while simultaneously being analyzed for the presence of solar activity to raise a trigger. On the arrival of a trigger the voltage buffer can be saved to disk and interferometric observations can be commenced with a denser subarray and using a larger observing bandwidth.
@@ 9 Acknowledgement
This scientific work uses data obtained from Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and Native Title Holders of the observatory site. Support for the operation of the MWA is provided by the Australian Government (NCRIS), under a contract to Curtin University administered by Astronomy Australia Limited. This work was supported by resources provided by the Pawsey Supercomputing Research Centre’s Setonix Supercomputer and Garrawarla GPU Cluster. Courtesy of National Institute of Information and Communications Technology (NICT). We thank the NICT for providing the Yamagawa data. We sincerely thanks Kazumasa Iwai (Nagoya University) for helping us to initiate conversion with NICT. We acknowledge Pietro Zucca for providing the LOFAR beam-formed data used in this work. D. P., S. D. and D. O. acknowledge the support of the Department of Atomic Energy, Government of India, under project no. 12-R&D-TFR-5.02-0700. D. K. acknowledges the support by the NASA Living with a Star Jack Eddy Postdoctoral Fellowship Program, administered by UCAR’s Cooperative Programs for the Advancement of Earth System Science (CPAESS) under award 80NSSC22M0097. We thank Barnali Das (CSIRO, Australia) for providing the name of the pipeline and the anomymous referees for constructive comments which improved the clarity and presentation of this work.

### v2

@@ 3.3 Background Subtraction
Here, $M_{T}[S^{\prime}(\nu=\nu_{0},t)]$ represents the moving median of pixel values from timestamp tTt-T to $t$ for frequency $\nu_{0}$.
@@
To get good estimates of the background noise, we stitch together two consecutive 30-minute data chunks and calculate the median by setting the window width, $T$, to 1800 s. This duration is significantly larger than the lifetime of most solar radio bursts at any given frequency, and has been chosen with the motivation of reducing the likelihood of any solar radio bursts present in the window impacting the median value.
Finally, this background-subtracted dynamic spectrum, $S_{\mathrm{bkgsub}}(\nu,t)$, is used for further analysis. The top panels of figure 7 show the dynamic spectrum before and after the median based background subtraction. As it can be clearly seen from the figure that the solar event has become much prominent.
@@ 4.1 Generating Binary Dynamic Spectrum
The temporal resolution of the data received from the Yamagawa is 1 second and hence, to avoid false detection due to the presence of any low-level transient RFI, we only search for events which are present for more than at least 3 s. For each spectral channel, we estimate the background levels by calculating the standard deviation ($\sigma(\nu)$) from the previous day quiet time dynamic spectrum. A signal-to-noise ratio (SNR) dynamic spectrum of the real-time data is created after the following operation on each spectral channel –
@@ 4.2 Event Detection Algorithm
We try to find closed contours on the image after morphological closing operation using OpenCV which uses border following method suzuki1985topological. Contours having minimum span of 15 MHz in frequency and 3 s in time are considered to be true events. This method not only finds closed contours but also filters out residual RFI which do not match the frequency and temporal span criteria. The bottom panels in figure 7 show the detection of such a closed contour region in the binary dynamic spectrum. If we find any such region within the most recent two minutes, we consider this event to be observed by triggering the MWA.
@@ 5 Triggering the MWA
The final step is to communicate with the MWA once a triggering condition is met and to schedule appropriate solar observations. This is achieved via a web-based triggering mechanism that has been developed for the MWA Hancock.
On arrival of a trigger, the buffer data is frozen and dumped to disk. As mentioned in Sec. 2.1, the buffer data is in ‘picket-fence’ mode. Simultaneously, an interferometric observation for the next 2 minutes is initiated. STORMY’s detection algorithm estimates the spectral span of the detected event. This information is used for defining the spectral coverage of the triggered observations. If $\nu_{min}$ and $\nu_{max}$ are the minimum and maximum detected frequency, the MWA observing bandwidth of 30.72 MHz is distributed linearly across $(\nu_{min}-30)$ MHz to $(\nu_{max}+10)$ MHz, subject to the MWA observing band of 8080300MHz300\ MHz. It also avoids the RFI afflicted channels in frequency range of 241–271 MHz while determining the channels for observation. The 30 MHz in lower limit and 10 MHz in upper limit are guard bands which allow for spectral drifts and the presence of weaker active emissions in the neighboring spectral regions which might not have been detectable at Yamagawa, but might be detectable by the more sensitive MWA. The 30MHz30\ MHz guard band on the lower frequency side is chosen to accommodate the typical drift rate of type II bursts of $0.1\ MHz\ s^{-1}$. Figure 8 shows how STORMY chooses the spectral channels for interferometric observations. In case of long duration bursts, as shown in the figure, STORMY continues to detect and adjust the frequency coverage for subsequent interferometric observations.
Figure 8: This figure shows how STORMY schedules subsequent MWAX observations after detection. The green rectangles are buffer channels. When STORMY triggers an observation (denoted by first blue vertical line), the 30.72 MHz band width is allocated according to the detected spectral span (red rectangles). In case of large bursts, as presented here, STORMY continues to adjust the frequency coverage as the burst drifts into lower frequency. The red transparent rectangle shows the RFI contaminated channels in MWA and are always excluded from observation.
@@ 7 Some Sample Events Captured by STORMY
    Figure 10: Top panels: (Left) Pre-processed Yamagawa dynamic spectrum. (Right) Radio image of the data captured at 126 MHz on 1 August 2024.The dotted white circle shows the optical disc of the Sun. The red contours are at the 1,10,30,50,90% of the peak. The ellipse inside the box at the bottom left corner shows the point-spread-function. Bottom panels: (Left and Middle) close-up of AIA images during the event. The dotted circles show the active region from which the eruption is taking place. (Right) Radio contours overlaid on the AIA map.
[…]
@@
8 Discussion and Conclusion
Though developed with the objective of increasing the observing efficiency of solar activity by the MWA, the framework of STORMY has wider applicability. Next, we discuss some of the few specific instances as examples where solar triggering can be implemented on modern radio interferometers:
  • LOFAR: The LOw Frequency ARrray is a versatile instrument, primarily due to its fully digital signal chain lofar_paper. It can observe multiple sources simultaneously via its multi-beam and subarray capabilities. The upcoming LOFAR 2.0 lofar2023whitepaper includes the LOFAR for Space Weather (LOFAR4SW) extension, which is planned to use few stations to track the Sun while it is above the instrument’s elevation limit commensal with other astronomical observations, simultaneously for the Low-Band Antennas (LBA; 10–90 MHz) and High-Band Antennas (HBA; 110–240 MHz). Using a part of the same instrument for solar monitoring offers several advantages over using a remote spectrograph - (a) the trigger comes from a monitoring system of better sensitivity and experiences RFI environment similar to the observing system, (b) co-location of the monitoring station removes the dependency on an independent distant system and eliminates the delays arising from the time taken to fetch remote data, substantially reducing the memory requirement of the ring buffer. Additionally, in LOFAR 2.0, having the ability to monitor the Sun simultaneously in HBA and LBA, can potentially enable one to raise an observing trigger for the LBA just a little before the emission drifts into the observing band of LBA. As an illustration, we demonstrate an application of STORMY to LOFAR data in figure 12. These data come from the LBA beam-formed mode, which was recorded with a spectral resolution of 12.207 kHz and temporal resolution of 10.486 ms, and have been averaged to 195.3 kHz and 1 s for faster processing. STORMY’s RFI excision and burst detection algorithms are seen to work very well on the LOFAR dynamic spectra. It took STORMY $\sim$30 $s$ to process the whole 40-minute datachunk. Hence, STORMY can be used on real-time beam-formed data to detect events and trigger interferometric observation when appropriate.
@@ 8 Discussion and Conclusion
  • SKAO: The triggering aspects are particularly relevant for the SKA-Low, which operates in the 50–350 MHz band where the bulk of solar radio activity is observed. Like the MWA, it uses aperture array elements and hence electronic pointing, essentially eliminating the slew time needed to point to the target source, post the trigger. The SKA-Low is expected to reach its AA* milestone before the end of this decade SKAO-YearInLife. It will have 307 stations, each comprising 256 dual-polarization dipoles log-periodic antennas placed over a footprint of 38m38\ m diameter. The SKA-Low supports flexible options for forming not only subarrays, but also substations, with the latter providing access to larger FoV as well as shorter baselines (within a station), though subject to bandwidth restrictions imposed by the correlator and the network capacity. AA* will support up to 1440 substations. The smallest substations have a 6m6\ m foot print and typically include 6 dual-pol antennas SKAO-Low-SubStation-Templates. A subarray comprising one such substation from $\sim$60 SKA-Low stations within a diameter of $\sim 10\ km$ will be an excellent match to the needs of triggered solar observing and is a referred to as solar subarray in the following text. It uses <0.5<0.5% of the SKA-Low collecting area, and hence has minimal sensitivity impact on the observations it is commensal with. Along the lines of the picket fence mode routinely employed at the MWA for solar observations, the solar subarray can use narrow spectral channels distributed across the 50–350 MHz SKA-Low observing band. For instance, using a hundred spectral channels, each 5.4 kHz in bandwidth, amounts to a mere 0.54 MHz of bandwidth. The solar subarray can be used to fill the voltage buffer while simultaneously being analyzed for the presence of solar activity to raise a trigger. SKA-Low has a substantial transient buffer, spanning 900 $s$ SKAO-YearInLife. To enable the capability of retrieving the event data lost due to time taken for trigger detection and generation, the solar subarray data can be directed to the buffer. On the arrival of a trigger the voltage buffer can be saved to disk and interferometric observations can be commenced with a denser subarray and using a larger observing bandwidth.
@@ 9 Acknowledgement
This scientific work uses data obtained from Inyarrimanha Ilgari Bundara, the CSIRO Murchison Radio-astronomy Observatory. We acknowledge the Wajarri Yamaji People as the Traditional Owners and Native Title Holders of the observatory site. Support for the operation of the MWA is provided by the Australian Government (NCRIS), under a contract to Curtin University administered by Astronomy Australia Limited. This work was supported by resources provided by the Pawsey Supercomputing Research Centre’s Setonix Supercomputer and Garrawarla GPU Cluster. We thank the National Institute of Information and Communications Technology (NICT) for providing the Yamagawa data. We sincerely thank Kazumasa Iwai (Nagoya University) for helping us to initiate conversation with NICT. We acknowledge Pietro Zucca for providing the LOFAR beam-formed data used in this work. D. P., S. D. and D. O. acknowledge the support of the Department of Atomic Energy, Government of India, under project no. 12-R&D-TFR-5.02-0700. D. K. acknowledges the support by the NASA Living with a Star Jack Eddy Postdoctoral Fellowship Program, administered by UCAR’s Cooperative Programs for the Advancement of Earth System Science (CPAESS) under award 80NSSC22M0097. We thank Barnali Das (CSIRO, Australia) for providing the name of the pipeline and the anomymous referees for constructive comments which improved the clarity and presentation of this work.